OpenStudy (anonymous):

medals rewarded! Eight soccer teams are to play each other two times in a season. How many games will be played? a.56 games b. b.32 games c.16 games d.40 games

OpenStudy (anonymous):

you can think of it this way: you are on a team, and your team plays 7 games and each other team plays 7 games so in total there are \(8\times 7=56\) but you have to divide by 2 because each game counts for two teams i hope b is 28

OpenStudy (anonymous):

if not, then i made a mistake

OpenStudy (anonymous):

28 isn't an answer choice

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

oh because i am not smart

OpenStudy (anonymous):

your team plays every other team TWICE so your team and each other team plays 14 games

OpenStudy (anonymous):

therefore it is \(\frac{8\times 14}{2}=8\times 7=56\)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw